How fast did the wagon move, with a mass of 20 t, if during the collision with the stop, each of its two systems was compressed by 10 cm. It is known that each spring is compressed 1 cm under the action of force 9800 N.


1
Expert's answer
2021-05-19T11:06:23-0400

According to the energy conservation law, the kinetic energy of the mooving wagon was converted to the potential energy of spring elastic deformation:


"\dfrac{mv^2}{2} =2\cdot \dfrac{kx^2}{2}"

where "m = 20t = 20\cdot 10^3kg" is the mass of the vagon, "v" is its initial velocity, "k" is the spring constant of each spring, "x = 10cm = 0.1m" is the compression rate of each spring. Factor 2 appeares since there are two identical springs in the system.

The spring constant can be found from the Hook's law:


"k = \dfrac{F}{x_0}"

where "F = 9800N" is the force, and "x_0 = 1cm = 0.01m" is the compression under this force.

Thus, obtain:


"v = x\sqrt{\dfrac{k}{m}} = x\sqrt{\dfrac{F}{x_0m}}\\
v = 0.1\cdot \sqrt{\dfrac{9800}{0.01\cdot 20\cdot 10^3}} = 0.7m/s"

Answer. 0.7 m/s.